# A Lie algebraic pattern behind logarithmic CFTs

### v2

@@ 1.1. Preliminary from Lie algebra
respectively. We sometimes identify the set of simple roots $\Pi=\{\alpha_{1},\dots,\alpha_{r}\}$ with the set of integers $I=\{1,\dots,r\}$. The fundamental weight $\varpi_{i}$ is the dual vector of the coroot. Denote $Q=\sum_{i\in I}\mathbb{Z}\alpha_{i}$, $Q^{\vee}=\sum_{i\in I}\mathbb{Z}\alpha_{i}^{\vee}$, $Q^{\ast}=\check{P}=\sum_{i\in I}\mathbb{Z}\alpha_{i}^{\ast}$ and $P=\sum_{i\in I}\mathbb{Z}\varpi_{i}$ the root lattice, coroot lattice, coweight lattice and weight lattice, respectively. The set of dominant integral weights $P_{+}$ and dominant integral coweights $\check{P}_{+}=Q^{\ast}_{+}$ are given by $P_{+}=\sum_{i\in I}\mathbb{Z}_{\geq 0}\varpi_{i}$, $\check{P}=\sum_{i\in I}\mathbb{Z}_{\geq 0}\alpha_{i}^{\ast}$, respectively. We use the letter $P_{\mathrm{min}}\subseteq P_{+}$ for the family of minuscule weights. Let $G$ be the corresponding simply-connected simple algebraic group with $B$ the Borel subgroup. Given a subset $J\subset I$, we denote by $P_{J}$ the corresponding parabolic subgroup [^5] Throughout the paper, we use the negative parabolic/Borel subalgebras. and $j\in J$, by $SL_{2}^{j}$ the corresponding subgroup isomorphic to $SL_{2}$. Denote $\rho=\sum_{i\in I}\varpi_{i}$ and $\rho^{\vee}=\sum_{i\in I}\alpha_{i}^{\ast}$ the Weyl vector and Weyl covector, $\theta$ and $\theta_{s}$ the highest (long) root ad highest short root, $h$ and $h^{\vee}$ the Coxeter and dual Coxeter number, respectively. For $\sigma\in W$, $l(\sigma)$ and $w_{0}$ denote the length of $\sigma$ and the longest element in $W$, respectively. For a minimal expression $\sigma=\sigma_{i_{n}}\cdots\sigma_{i_{1}}$ of $\sigma\in W$, we sometimes use $\sigma_{i_{0}}=\operatorname{id}$ for convenience. For $\beta\in P_{+}$, denote $L_{\beta}$ the finite-dimensional irreducible $\mathfrak{g}$-module with the highest weight $\beta$. We sometimes use the letter $y_{\beta}$ and $y^{\prime}_{\beta}$ for a highest weight vector and lowest weight vector of $L_{\beta}$, respectively. The Langlands dual ${}^{L}\mathfrak{g}$ is defined by replacing the root system by the coroot system. Namely,
@@
(in particular, $\rho_{{}^{L}\mathfrak{g}}=\rho^{\vee}$ and $\rho_{{}^{L}\mathfrak{g}}^{\vee}=r^{\vee}\rho$). If we want to emphasize that a certain object $X$ (e.g., $\mathfrak{g}$, $Q$,…) is derived from ${}^{L}\mathfrak{g}$ rather than $\mathfrak{g}$, we often use symbols such as ${}^{L}X$, $X_{{}^{L}\mathfrak{g}}$, $\check{X}$, etc. By abuse of notation, for $\theta=\sum_{i\in I}a_{i}\alpha_{i}$, denote ${}^{L}\theta=\sum_{i\in I}a_{i}\alpha_{r+1-i}^{\vee}\in\mathfrak{h}^{\ast}$ (i.e., we regard ${}^{L}\theta_{\mathfrak{g}}=\theta_{{}^{L}\mathfrak{g}}$ as an element in $\mathfrak{h}^{\ast}$). For more detailed data in each cases, see e.g. [Kac90, p.91-92] (but note that the labeling of Dynkin diagrams are different).
@@
[Sug21, (114), (115), Lemma 4.5] For a $P_{i}$-module $M$ and $\mu\in P$, we have
@@
It suffices to show $\mathrm{ad}_{e_{i}}^{1-c_{ij}}(e_{j})=0$ for $i\neq j$ on $\bigcap_{j\in J}M_{j}$. Since $[e_{i},f_{j}]=0$, we have
@@ 1.2. Shift system
In this subsection, we introduce a new concept, named shift system, which is the main tool in the present paper, and give its basic examples and properties.
@@
      For any $i\in\Pi$ and $\lambda\not\in\Lambda^{\sigma_{i}}$, there exists $P_{i}$-submodules $W_{i,\lambda}\subseteq V_{\lambda}$, $W_{i,\sigma_{i}\ast\lambda}\subseteq V_{\sigma_{i}\ast\lambda}$ and a $B$-module homomorphism $Q_{i,\lambda}\colon V_{\lambda}\rightarrow W_{i,\sigma_{i}\ast\lambda}({\sigma_{i}}\uparrow{\lambda})$ such that we have the short exact sequence
@@
For convenience, let us introduce some terms. For a given shift system, we call $V_{\lambda}$ a staggered $B$-module. The action of $f_{i}\in\mathfrak{n}_{-}$ on $V_{\lambda}$ and the $B$-module homomorphism $Q_{i,\lambda}$ above are referred to as a long screening operator and a short screening operator, respectively. If some object $X$ is isomorphic (resp. conjectured to be isomorphic) to $H^{0}(P\times_{B}V_{\lambda})$ for a staggered $B$-module $V_{\lambda}$ and a parabolic subgroup $P\supseteq B$, we call $H^{0}(P\times_{B}V_{\lambda})$ the Feigin-Tipunin (FT) construction of $X$ (resp. FT conjecture on $X$).
@@
    By substituting $\sigma=\operatorname{id}$ to the axiom (2a), we have $\operatorname{id}\uparrow\lambda=0$. On the other hand, substituting $\sigma=\sigma_{i}$ shows that $\sigma_{i}\uparrow\lambda+\sigma_{i}\uparrow(\sigma_{i}\ast\lambda)=-\alpha_{i}$ ($\lambda\not\in\Lambda^{\sigma_{i}}$) or $-2\alpha_{i}$ ($\lambda\in\Lambda^{\sigma_{i}}$). By axioms (2a) and (2b), if $l(\sigma_{i}\sigma)=l(\sigma)-1$, then $(\sigma\uparrow\lambda,\alpha_{i}^{\vee})\in\mathbb{Z}_{<0}$.
@@
    For any $1\leq m\leq n$, $(\sigma_{i_{m}}\cdots\sigma_{i_{1}})\uparrow\lambda=\sum_{j=1}^{m}\sigma_{i_{j}}\uparrow(\sigma_{i_{j-1}}\cdots\sigma_{i_{0}}\ast\lambda)$. (1.2)
@@
    For $j\in\Pi$, $\beta\in P$ such that $(\beta,\alpha_{j}^{\vee})\geq 0$ and $v\in V_{\lambda}^{h=\beta}$, we have $v=0$ (resp. $v\in W_{j,\lambda}$) iff $f_{j}^{(\beta,\alpha_{j}^{\vee})}v=0$ (resp. $f_{j}^{(\beta,\alpha_{j}^{\vee})+1}v=0$).
@@
$H^{0}(G\times_{B}V_{\lambda})\simeq\bigoplus_{\alpha\in P_{+}\cap Q}L_{\beta}\otimes\mathcal{W}_{-\alpha+\lambda}.$ (1.3)
In Part II, we will consider the cases where $\Lambda$ is the unique representative of $\tfrac{1}{p}Q^{\ast}/Q$ for the $W$-action defined by $x=\tfrac{1}{p}\rho^{\vee}$ ($p=r^{\vee}m$, $m\in\mathbb{Z}_{\geq 1}$) and $x=\tfrac{1}{p}\rho$ ($p=2m-1$, $m\in\mathbb{Z}_{\geq 1}$ and $\mathfrak{g}=B_{r}$), respectively.
The following lemma is shown in exactly the same manner as [Sug21], so the proof is omitted.
@@
[Sug21, Lemma 3.17-3.18] Let $V_{\lambda}$ be a staggered $B$-module and $x\in V_{\lambda}^{h=\beta}$ a nonzero vector for some $\beta\in P_{+}$. If there exists a $B$-module homomorphism $\Phi\colon L_{\beta}\rightarrow U(\mathfrak{b})x$ that sends the highest weight vector $y_{\beta}$ of $L_{\beta}$ to $x$, then $\Phi$ is an isomorphism. In particular, $U(\mathfrak{b})x$ is a $G$-submodule of $V_{\lambda}$.
1.3. Feigin-Tipunin conjecture/construction
Let $(\Lambda,\uparrow,\{V_{\lambda}\}_{\lambda\in\Lambda})$ be a shift system. In this subsection, we will give a brief necessary and sufficient condition for the Feigin-Tipunin conjecture [FT10, Sug21]
@@ 1.3. Feigin-Tipunin conjecture/construction
to hold. For $(i,j,\lambda)\in I\times I\times\Lambda$, we consider the following condition (see [Sug21, (98),(99),(100)]):
[…]
@@
$(i,j,\lambda)$ satisfies (1.4) for any $(i,j)\in I\times I$. (1.6)
@@
[Sug21, Lemma 4.3] Let $i,j\in I$ and $\lambda\in\Lambda$. If $(i,j,\lambda)$ satisfies (1.4), then $W_{i,\lambda}\cap W_{j,\lambda}$ is closed under the operator $e_{i}$. In particular, if $(J,\lambda)\subseteq I\times\Lambda$ satisfies (1.5), then $\bigcap_{j\in J}W_{j,\lambda}$ is a $P_{J}$-module.
@@
$f_{i}^{(\beta,\alpha_{i}^{\vee})}Q_{j,\lambda}A_{\beta}=Q_{j,\lambda}f_{i}^{(\beta,\alpha_{i}^{\vee})}A_{\beta}=0,$ (1.7)
@@
[Sug21, Theorem 4.4] Only if $(J,\lambda)$ satisfies (1.5), then $\bigcap_{j\in J}W_{j,\lambda}$ is a $P_{J}$-submodule of $V_{\lambda}$.
@@
$0\not=f_{i}^{(\beta,\alpha_{i}^{\vee})}w\in\bigcap_{i\not=k\in J}(W_{k,\lambda}^{h=\sigma(\beta)+\sigma_{i}\uparrow\lambda})^{e_{k}}.$ (1.8)
@@
Since the proof is the same as [Sug21, Lemma 4.15 - 4.18], we omit the detail. For $\beta\in P_{+}$, denote $W_{\beta}$ the set of highest weight vectors of $H^{0}(G\times_{B}V_{\lambda})$ with highest weight $\beta$. Let $G/B=\bigcap_{\sigma\in W}U_{\sigma}$ be the Schubert open covering. For $s\in W_{\beta}$, we have $s(N_{+})=s(\operatorname{id})$, and thus $s|_{U_{\operatorname{id}}}=1\otimes v$ for some $v\in V_{\lambda}^{h=\beta}$. By considering coordinate changes between $U_{\sigma}$’s, $\operatorname{ev}|_{W_{\beta}}$ is injective. By Lemma 1.9, $\operatorname{ev}|_{L_{\beta}\otimes W_{\beta}}$ so is. Now we just need to make sure that $L_{\beta}\otimes W_{\beta}$’s are linearly independent with respect to different highest weights $\beta$’s. For a fixed $\gamma\in P$ and $\Delta$, let us take a vector $s=\sum_{\beta\in P_{+}}s_{\beta}\in H^{0}(G\times_{B}V_{\lambda,\Delta})^{h=\gamma}$ such that $\operatorname{ev}(s)=0$. We fix a minimal highest weight $\alpha\in P_{+}$ in $\{\beta\in P_{+}\ |\ s_{\beta}\not=0\}$ with respect to the standard partial order $\geq$ on $\mathfrak{h}^{\ast}$, i.e. $\mu\geq\mu^{\prime}$ iff $\mu-\mu^{\prime}\in\bigoplus_{i\in I}\mathbb{Z}_{\geq 0}\alpha_{i}$. Since $L_{\beta_{2}}^{h=w_{0}(\beta_{1})}=\{0\}$ for $\beta_{1}\not\leq\beta_{2}$, there exists $X\in U(\mathfrak{b})$ such that
@@
Let us prove Theorem 1.1(1) (see also [Sug21, Section 4.5]). By Theorem 1.12 and Remark 1.7(3), the first half of Theorem 1.1(1) is proved. By Lemma 1.10 and Lemma 1.11, $\bigcap_{i\in I}W_{i,\lambda}$ has the $G$-module structure (and thus, in $H^{0}(G\times_{B}V_{\lambda})$) if and only if $\lambda\in\Lambda$ satisfies the weak condition (weak). This completes the proof.
1.4. Borel-Weil-Bott type theorem

### v3

@@ 1.1. Preliminary from Lie algebra
respectively. We sometimes identify the set of simple roots $\Pi=\{\alpha_{1},\dots,\alpha_{r}\}$ with the set $I=\{1,\dots,r\}$. The fundamental weight $\varpi_{i}$ is the dual vector of the coroot. Denote $Q=\sum_{i\in I}\mathbb{Z}\alpha_{i}$, $Q^{\vee}=\sum_{i\in I}\mathbb{Z}\alpha_{i}^{\vee}$, $Q^{\ast}=\check{P}=\sum_{i\in I}\mathbb{Z}\alpha_{i}^{\ast}$ and $P=\sum_{i\in I}\mathbb{Z}\varpi_{i}$ the root lattice, coroot lattice, coweight lattice and weight lattice, respectively. The set of dominant integral weights $P_{+}$ and dominant integral coweights $\check{P}_{+}=Q^{\ast}_{+}$ are given by $P_{+}=\sum_{i\in I}\mathbb{Z}_{\geq 0}\varpi_{i}$, $\check{P}=\sum_{i\in I}\mathbb{Z}_{\geq 0}\alpha_{i}^{\ast}$, respectively. We use the letter $P_{\mathrm{min}}\subseteq P_{+}$ for the family of minuscule weights. Let $G$ be the corresponding simply-connected simple algebraic group with $B$ the Borel subgroup. Given a subset $J\subset I$, we denote by $P_{J}$ the corresponding parabolic subgroup [^5] Throughout the paper, we use the negative parabolic/Borel subalgebras. and $j\in J$, by $SL_{2}^{j}$ the corresponding subgroup isomorphic to $SL_{2}$. Denote $\rho=\sum_{i\in I}\varpi_{i}$ and $\rho^{\vee}=\sum_{i\in I}\alpha_{i}^{\ast}$ the Weyl vector and Weyl covector, $\theta$ and $\theta_{s}$ the highest (long) root and highest short root, $h$ and $h^{\vee}$ the Coxeter and dual Coxeter number, respectively. For $\sigma\in W$, $l(\sigma)$ and $w_{0}$ denote the length of $\sigma$ and the longest element in $W$, respectively. For a minimal expression $\sigma=\sigma_{i_{n}}\cdots\sigma_{i_{1}}$ of $\sigma\in W$, we sometimes use $\sigma_{i_{0}}=\operatorname{id}$ for convenience. For $\beta\in P_{+}$, denote $L_{\beta}$ the finite-dimensional irreducible $\mathfrak{g}$-module with the highest weight $\beta$. We sometimes use the letter $y_{\beta}$ and $y^{\prime}_{\beta}$ for a highest weight vector and lowest weight vector of $L_{\beta}$, respectively. The Langlands dual $^{L}\mathfrak{g}$ is defined by replacing the root system by the coroot system. Namely,
@@
(in particular, $\rho_{{}^{L}\mathfrak{g}}=\rho^{\vee}$ and $\rho_{{}^{L}\mathfrak{g}}^{\vee}=r^{\vee}\rho$). If we want to emphasize that a certain object $X$ (e.g., $\mathfrak{g}$, QQ,…) is derived from $^{L}\mathfrak{g}$ rather than $\mathfrak{g}$, we often use symbols such as $^{L}X$, $_{{}^{L}\mathfrak{g}}, Xˇ\check{X}$, etc. By abuse of notation, for $\theta=\sum_{i\in I}a_{i}\alpha_{i}$, denote $^{L}\theta=\sum_{i\in I}a_{i}\alpha_{r+1-i}^{\vee}\in\mathfrak{h}^{\ast}$ (i.e., we regard $^{L}\theta_{\mathfrak{g}}=\theta_{{}^{L}\mathfrak{g}}$ as an element in $\mathfrak{h}^{\ast}$). For more detailed data in each cases, see e.g. [34, p.91-92] (but note that the labeling of Dynkin diagrams are different).
@@
[45, (114), (115), Lemma 4.5] For a $P_{i}$-module $M$ and $\mu\in P$, we have
@@
It suffices to show $\mathrm{ad}_{e_{i}}^{1-c_{ij}}(e_{j})=0$ for $i\neq j$ on $\bigcap_{j\in J}M_{j}$. Since $e_{i},f_{j}]=0$, we have
@@ 1.2. Shift system
In this subsection, we introduce a new concept, named shift system [^6] The name “shift system” was chosen because, as we show below, the shift $\sigma\uparrow\lambda$ plays a central role in our representation theory. However, this is tentative (an alternative name could be FT data, Felder/screening system, etc.) and may be changed to a more appropriate name/concept as our understanding of the theory of FT construction improves., which is the main tool in the present paper, and give its basic examples and properties.
@@
      For any $i\in I$ and $\lambda\not\in\Lambda^{\sigma_{i}}$, there exists $P_{i}$-submodules $W_{i,\lambda}\subseteq V_{\lambda}$, $W_{i,\sigma_{i}\ast\lambda}\subseteq V_{\sigma_{i}\ast\lambda}$ and a $B$-module homomorphism $Q_{i,\lambda}\colon V_{\lambda}\rightarrow W_{i,\sigma_{i}\ast\lambda}({\sigma_{i}}\uparrow{\lambda})$ such that we have the short exact sequence
@@
Let us introduce some terms for convenience. The action of $f_{i}\in\mathfrak{n}_{-}$ on $V_{\lambda}$ and the $B$-module homomorphism $Q_{i,\lambda}$ above are referred to as a long screening operator and a short screening operator, respectively. If some object $X$ is isomorphic (resp. conjectured to be isomorphic) to $H^{0}(P\times_{B}V_{\lambda})$ for some $V_{\lambda}$ and a parabolic subgroup $P\supseteq B$, we call $H^{0}(P\times_{B}V_{\lambda})$ the Feigin–Tipunin construction of $X$ (resp. FT conjecture on $X$).
@@
    By substituting $\sigma=\operatorname{id}$ to the axiom (2a), we have $\operatorname{id}\uparrow\lambda=0$. On the other hand, substituting $\sigma=\sigma_{i}$ shows that $\sigma_{i}\uparrow\lambda+\sigma_{i}\uparrow(\sigma_{i}\ast\lambda)=-\alpha_{i}$ ($\lambda\not\in\Lambda^{\sigma_{i}}$) or $2\alpha_{i}$ ($\lambda\in\Lambda^{\sigma_{i}}$). By axioms (2a) and (2b), if $l(\sigma_{i}\sigma)=l(\sigma)-1$, then $(\sigma\uparrow\lambda,\alpha_{i}^{\vee})\in\mathbb{Z}_{<0}$.
@@
    (1.3) For any $1\leq m\leq n$, $(\sigma_{i_{m}}\cdots\sigma_{i_{1}})\uparrow\lambda=\sum_{j=1}^{m}\sigma_{i_{j}}\uparrow(\sigma_{i_{j-1}}\cdots\sigma_{i_{0}}\ast\lambda)$.
@@
    For $j\in I$, $\beta\in P$ such that $(\beta,\alpha_{j}^{\vee})\geq 0$ and $v\in V_{\lambda}^{h=\beta}$, we have $v=0$ (resp. $v\in W_{j,\lambda}$) iff $f_{j}^{(\beta,\alpha_{j}^{\vee})}v=0$ (resp. $f_{j}^{(\beta,\alpha_{j}^{\vee})+1}v=0$).
@@
(1.4) $H^{0}(G\times_{B}V_{\lambda})\simeq\bigoplus_{\alpha\in P_{+}\cap Q}L_{\alpha+\lambda^{\bullet}}\otimes\mathcal{W}_{-\alpha+\lambda}.$
In Part II, we will consider the cases where $\Lambda$ is the unique representative of $\tfrac{1}{p}Q^{\ast}/Q$ for the $W$-action defined by $x=\tfrac{1}{p}\rho^{\vee}$ ($p=r^{\vee}m$, $m\in\mathbb{Z}_{\geq 1}$) and $x=\tfrac{1}{p}\rho$ (p=2m1p=2m-1, $m\in\mathbb{Z}_{\geq 1}$ and $\mathfrak{g}=B_{r}$), respectively.
The following lemma is shown in exactly the same manner as [45], so the proof is omitted.
@@
[45, Lemma 3.17-3.18] Let $x\in V_{\lambda}^{h=\beta}$ be a nonzero vector for some $\beta\in P_{+}$. If there exists a $B$-module homomorphism $\Phi\colon L_{\beta}\rightarrow U(\mathfrak{b})x$ that sends the highest weight vector $y_{\beta}$ of $L_{\beta}$ to $x$, then $\Phi$ is an isomorphism. In particular, $U(\mathfrak{b})x$ is a $G$-submodule of $V_{\lambda}$.
1.3. Feigin–Tipunin conjecture/construction
Let $(\Lambda,\uparrow,\{V_{\lambda}\}_{\lambda\in\Lambda})$ be a shift system. In this subsection, we will give a brief necessary and sufficient condition for the Feigin-Tipunin conjecture [27, 45]
@@ 1.3. Feigin–Tipunin conjecture/construction
to hold. For $(i,j,\lambda)\in I\times I\times\Lambda$, we consider the following condition (see [45, (98),(99),(100)]):
(1.5) $\text{$\lambda\in\Lambda^{\sigma_{j}}$ or $({\sigma_{j}}\uparrow{\lambda},\alpha_{i}^{\vee})=-\delta_{ij}$}.$
@@
(1.7) $(i,j,\lambda)$ satisfies (1.5) for any $(i,j)\in I\times I$.
@@
[45, Lemma 4.3] Let $i,j\in I$ and $\lambda\in\Lambda$. If $(i,j,\lambda)$ satisfies (1.5), then $W_{i,\lambda}\cap W_{j,\lambda}$ is closed under the operator $e_{i}$. In particular, if $(J,\lambda)\subseteq I\times\Lambda$ satisfies (1.6), then $\bigcap_{j\in J}W_{j,\lambda}$ is a $P_{J}$-module.
@@
(1.8) $f_{i}^{(\beta,\alpha_{i}^{\vee})}Q_{j,\lambda}A_{\beta}=Q_{j,\lambda}f_{i}^{(\beta,\alpha_{i}^{\vee})}A_{\beta}=0,$
@@
[45, Theorem 4.4] Only if $(J,\lambda)$ satisfies (1.6), then $\bigcap_{j\in J}W_{j,\lambda}$ is a $P_{J}$-submodule of $V_{\lambda}$.
@@
(1.9) $0\not=f_{i}^{(\beta,\alpha_{i}^{\vee})}w\in\bigcap_{i\not=k\in J}(W_{k,\lambda}^{h=\sigma(\beta)+\sigma_{i}\uparrow\lambda})^{e_{k}}.$
@@
Since the proof is the same as [45, Lemma 4.15 - 4.18], we omit the detail. For $\beta\in P_{+}$, denote $W_{\beta}$ the set of highest weight vectors of $H^{0}(G\times_{B}V_{\lambda})$ with highest weight $\beta$. Let $G/B=\bigcap_{\sigma\in W}U_{\sigma}$ be the Schubert open covering. For $s\in W_{\beta}$, we have $s(N_{+})=s(\operatorname{id})$, and thus $s|_{U_{\operatorname{id}}}=1\otimes v$ for some $v\in V_{\lambda}^{h=\beta}$. By considering coordinate changes between $U_{\sigma}$’s, $\operatorname{ev}|_{W_{\beta}}$ is injective. By Lemma 1.9, $\operatorname{ev}|_{L_{\beta}\otimes W_{\beta}}$ so is. Now we just need to make sure that $L_{\beta}\otimes W_{\beta}$’s are linearly independent with respect to different highest weights $\beta$’s. For a fixed $\gamma\in P$ and $\Delta$, let us take a vector $s=\sum_{\beta\in P_{+}}s_{\beta}\in H^{0}(G\times_{B}V_{\lambda,\Delta})^{h=\gamma}$ such that $\operatorname{ev}(s)=0$. We fix a minimal highest weight $\alpha\in P_{+}$ in $\{\beta\in P_{+}\ |\ s_{\beta}\not=0\}$ with respect to the standard partial order $\geq$ on $\mathfrak{h}^{\ast}$, i.e. $\mu\geq\mu^{\prime}$ iff $\mu-\mu^{\prime}\in\bigoplus_{i\in I}\mathbb{Z}_{\geq 0}\alpha_{i}$. Since $L_{\beta_{2}}^{h=w_{0}(\beta_{1})}=\{0\}$ for $\beta_{1}\not\leq\beta_{2}$, there exists $X\in U(\mathfrak{b})$ such that
@@
Let us prove Theorem 1.1(1) (see also [45, Section 4.5]). By Theorem 1.12 and Remark 1.7(3), the first half of Theorem 1.1(1) is proved. By Lemma 1.10 and Lemma 1.11, $\bigcap_{i\in I}W_{i,\lambda}$ has the $G$-module structure (and thus, in $H^{0}(G\times_{B}V_{\lambda})$) if and only if $\lambda\in\Lambda$ satisfies the weak condition (weak). This completes the proof.
1.4. Borel–Weil–Bott-type theorem